# devoted to memory of Alexei Alexandrovich Starobinsky

### v12

@@ II.7 Hubble volume vs the volume of Universe, and thermal fluctuations of de Sitter state
The vacuum compressibility is not constrained by the thermodynamic laws and is determined by the ultraviolet physicsKlinkhamerVolovik2008 with its Planck energy scale, $\chi_{\rm vac}^{-1}\sim M_{\rm Pl}^{4}$. In this relation it is similar to the gravitational coupling, which is also determined by the Planck scale, $K\sim M_{\rm Pl}^{2}$. On the other hand, the temperature corrections to $\chi_{\rm vac}^{-1}$ and $K$ as well as the Casimir corrections are within the range of the infrared physics, see Ref. VolovikZelnikov2003 for the universal temperature correction to the gravitational coupling $K$. The contribution to the compressibility from infrared range was discussed in Refs.BLHu2022 ; BLHu2024 ; Widom1998 . The negative contributions to compressibility obtained in these papers do not violate the stability of the quantum vacuum, since these contributions represent the corrections, which are small compared to the main value of the vacuum compressibility, $\chi_{\rm vac}^{-1}\sim M_{\rm Pl}^{4}$.
@@
III Thermodynamics from the heat transfer in the multi-metric gravity ensemble
III.1 Multi-metric gravity
@@
${\cal L}_{n}=e_{abcd}\left(K_{n}\,R^{ab(n)}\wedge e^{c(n)}\wedge e^{d(n)}+\Lambda_{n}\,e^{a(n)}\wedge e^{b(n)}\wedge e^{c(n)}\wedge e^{d(n)}\right)\,,$ (14)
@@
III.2 Heat exchange in multi-metric gravity
@@
III.3 Thermodynamics from the multi-metric ensemble
@@
The physical temperature in turn gives rise to the total entropy and to the local entropy, $S_{V}=\sum_{n}S_{n}=12\pi KHV=s_{\rm loc}V$. So, in this scenario the de Sitter entropy comes from a set of many randomly distributed subsystems with the expansion rates $H_{n}$. Due to the heat exchange at fixed total energy $E_{V}$ these states are organized in the equilibrium thermal state, which corresponds to the coherent de Sitter expansion of the whole system.
III.4 Regularization vs thermalization
@@
One may expect that the heat exchange between subsystems leads not only to the coherence of the de Sitter states, but to the general coherence of the metric fields, when the metric fields $g_{\mu\nu(n)}$ of the subsystems become equal, thus forming the common metric $g_{\mu\nu}$. If this is true, this could be some kind of the thermodynamic gravity, but without the holographic principle.
III.5 Coherence vs thermalization
@@
At first glance this formation of the time dependent coherent state looks similar to the formation of the coherent spin precession (magnon BEC) from the incoherent precessions about the local magnetic fields with random frequencies $\omega_{n}$.BunkovVolovik2013 The latter corresponds to random $H_{n}$. The coherence develops due to the spin currents between the regions of local precessions – the analog of the heat exchange. The common frequency $\omega$ of the coherent precession corresponds to the common Hubble parameter $H$. They have the same dimension, $[H]=[\omega]=[1/t]$, while the dimensionless magnon number (or the spin projection $S_{z}$) corresponds to the entropy which is also dimensionless. However, the source of the coherence is the exchange of spin instead of exchange of the energies, and also the formation of coherent state is due to minimization of the total energy $E$ at fixed projection $S_{z}$ of the total spin on magnetic field. This leads to the common frequency, which plays the role of the chemical potential $\mu$, which emerges due to the exchange of particles between the regions. That is why this process is quite opposite to the formation of temperature and entropy, where energy is fixed, while entropy reaches maximum value. This process corresponds to the minimization of energy at fixed total entropy. Anyway, in all the cases the parameters, which are the same in all subsystems in equilibrium – common frequency $\omega$, common chemical potential $\mu$, common temperature $T$ and common Hubble parameter $H$ – have the same dimension, $[T]=[H]=[\mu]=[\omega]=[1/t]$.
IV Contracting de Sitter and black hole entropy
IV.1 de Sitter vs black hole
@@ IV Contracting de Sitter and black hole entropy
IV.2 Entropy of expanding, contracting and static de Sitter
The static de Sitter vacuum with singularity at the cosmological horizon is the intermediate state between the two opposite states – the expanding and contracting states. The entropy of the static state is zero, and correspondingly the temperature must be infinite, see also Ref.Milekhin2024 . In the contracting de Sitter vacuum with $H<0$ the local temperature is negative, $T=H/\pi<0$, and the local entropy is also negative, $s_{\rm contracting}=3H/4G=-3/(4Gr_{h})<0$. The total entropy in the Hubble volume of contracting de Sitter is
@@ IV.2 Entropy of expanding, contracting and static de Sitter
where $A$ is the area of the cosmological horizon.
IV.3 Gravastar – black hole with de Sitter core
The black hole can be formed from the gravastar. This is the object, which contains the de Sitter state inside the black hole horizon.Chapline2003 ; Mazur2023 ; Mottola2023 We consider the gravastar in which the black hole horizon coincides with the de Sitter horizon, $r_{\rm bh}=r_{h}$ (the metric in the state with the critical value of the mass parameter $m=2MG|H|=1$ at which two horizons merge is illustrated in Fig. 1 of Ref. Dymnikova2002 ). Such gravastar is stationary, it has no Hawking radiation and its entropy is zero. In the Painlevé-Gullstrand form the metric is given by Eq.(1) with the following shift velocity:
@@ IV.3 Gravastar – black hole with de Sitter core
IV.4 Entropy of black hole from negative entropy of contracting de Sitter
According to Eq.(27) the region of the contracting de Sitter state has negative entropy, $S_{\rm contracting}=-A/4G$. That is why the gravastar is unstable towards the shrinking of the de Sitter region, since this leads to the increase of the entropy of the whole system due to decrease of the negative entropy of the contracting de Sitter state. Due to the energy conservation the singularity is formed ar $r=0$. In the final state – the black hole – the de Sitter region with negative entropy fully disappears by shrinking to the singularity with mass $M$. The resulting black hole with mass $M$ acquires the positive entropy, $A/4G$:
@@ IV.4 Entropy of black hole from negative entropy of contracting de Sitter
In terms of the entropies of horizons, this means that in the initial gravastar state the entropy of the contracting de Sitter horizon fully compensates the entropy of the black hole horizon:
@@
IV.5 White hole and anti-gravastar
@@
It is obtained from the pure white hole, which has the negative horizon entropy $S_{\rm WH}=-A/4G$,Volovik2022 ; Volovik2020 by growing the de Sitter state in its core with positive entropy. In the anti-gravastar state the entropies of two horizons cancel each other:
@@
IV.6 Heat exchange between black holes in the multi-metric ensemble
@@
V Thermodynamics of de Sitter state and $f(R)$ gravity
V.1 Thermodynamic variables in $f(R)$ gravity
@@
V.2 Gibbs-Duhem relation in $f(R)$ gravity
@@
V.3 Entropy of cosmological horizon in terms of effective gravitational coupling
@@
V.4 Example of quadratic gravity
@@
VI From de Sitter thermodynamics to de Sitter decay
VI.1 de Sitter state as thermal bath for matter
@@ VI From de Sitter thermodynamics to de Sitter decay
The nonzero local temperature of the gravitational vacuum suggests that the de Sitter vacuum is locally unstable towards the creation of thermal matter from the vacuum by thermal activation. This is distinct from the mechanism of creation of the pairs of particles by Hawking radiation from the cosmological horizon, which may or may not lead to the decay of the vacuum energy. There are still controversies concerning the stability of the de Sitter vacuum caused by Hawking radiation, see e.g. Kamenshchik2022 ; Polyakov2012 ; Polyakov2022 and references therein.
VI.2 de Sitter decay due to thermalization of matter by de Sitter heat bath
To describe the decay of the vacuum due to activation and thermalization of matter, the extension of the Starobinsky analysis of the vacuum decayStarobinsky1994 ; Starobinsky1996 ; Starobinsky1997 ; Starobinsky2023 is needed. Especially the revolutionary stochastic inflation approach pioneered by StarobinskyStarobinsky1982 ; Starobinsky1986 is extremely useful, although it requires some modifications.Pattison2019 ; Cruces2022 ; Cruces2022b ; Shellard2024 This also includes the so-called separate universe approach, which is somewhat similar to the multi-metric gravity discussed in Sec. III. Here we consider the simple phenomenological scenario based on the energy exchange between the thermal de Sitter and the thermal matter.
@@
Here we use for simplicity the conventional general relativity with $\epsilon_{\rm vac}=\Lambda=-\frac{1}{2}K_{0}{\cal R}$. This phenomenological description of the energy exchange between the de Sitter vacuum and matter does not depend on the details of the microscopic (UV) theory, and requires only the slow-roll condition, i.e. the slow variation of the Hubble parameter, $|\dot{H}|\ll H^{2}$, or which is the same $|\dot{T}|\ll T^{2}$.
Since the vacuum energy density is $\epsilon_{\rm vac}\propto KH^{2}$, one obtains from Eq.(52) the following time dependence of the Hubble parameter and of energy densities:
$H\sim M_{\rm Pl}\left(\frac{t_{\rm Pl}}{t+t_{0}}\right)^{1/3}\,,$ (53)
$\epsilon_{\rm vac}\sim M_{\rm Pl}^{4}\left(\frac{t_{\rm Pl}}{t+t_{0}}\right)^{2/3}\,,$ (54)
$\epsilon_{M}=bH^{4}\sim M_{\rm Pl}^{4}\left(\frac{t_{\rm Pl}}{t+t_{0}}\right)^{4/3}\,.$ (55)
@@
$H(t=0)\sim M_{\rm Pl}\left(\frac{t_{\rm Pl}}{t_{0}}\right)^{1/3}\ll M_{\rm Pl}\,.$ (56)
@@
So, the scenario of thermalization of matter by the de Sitter heat bath gives Eqs. (51) and (52). They produce the inflation in terms of two phenomenological parameters, $b$ and $t_{0}$, which determine the decay of the vacuum energy density:
$\epsilon_{\rm vac}(t)=6M_{\rm Pl}^{4}\left(\frac{t_{\rm Pl}}{b(t+t_{0})}\right)^{2/3}\,.$ (57)
VI.3 de Sitter decay and Zel’dovich stiff matter
As was mentioned by Padmanabhan,Padmanabhan2005 his model leads to a late time cosmological constant in Eq.(57), which is independent of the initial value, but its value is still far too large. Can we fix this? In Sec. V.2 we obtained that the thermodynamics of de Sitter thermal bath corresponds to the Zel’dovich stiff matter with $w=1$. Let us try the stiff matter scenario using the phenomenological approach.

### v15

@@ III.1 Thermal fluctuations of de Sitter state
III.2 Vacuum compressibility
The vacuum compressibility is not constrained by the thermodynamic laws and is determined by the ultraviolet physicsKlinkhamerVolovik2008 with its Planck energy scale, $\chi_{\rm vac}^{-1}\sim M_{\rm Pl}^{4}$. In this relation it is similar to the gravitational coupling, which is also determined by the Planck scale, $K\sim M_{\rm Pl}^{2}$. On the other hand, the temperature corrections to $\chi_{\rm vac}^{-1}$ and $K$ as well as the Casimir corrections are within the range of the infrared physics, see Ref. VolovikZelnikov2003 for the universal temperature correction to the gravitational coupling $K$. The contribution to the compressibility from the infrared range was discussed in Refs.BLHu2022 ; BLHu2024 ; BLHu2024a ; Widom1998 . The negative contributions to compressibility obtained in these papers do not violate the stability of the quantum vacuum, since these contributions represent the corrections, which are small compared to the main value of the vacuum compressibility, $\chi_{\rm vac}^{-1}\sim M_{\rm Pl}^{4}$.
@@ III.2 Vacuum compressibility
IV Thermodynamics from the heat transfer in the multi-metric gravity ensemble
IV.1 Multi-metric gravity
@@
${\cal L}_{n}=e_{abcd}K_{n}\,R^{ab(n)}\wedge e^{c(n)}\wedge e^{d(n)}+$
$+e_{abcd}\Lambda_{n}\,e^{a(n)}\wedge e^{b(n)}\wedge e^{c(n)}\wedge e^{d(n)}\,,$ (22)
@@
IV.2 Heat exchange in multi-metric gravity
@@
IV.3 Thermodynamics from the multi-metric ensemble
@@
The physical temperature in turn gives rise to the total entropy and to the local entropy, $S_{V}=\sum_{n}S_{n}=12\pi KHV=s_{\rm loc}V$. So, in this scenario the de Sitter entropy comes from a set of many randomly distributed subsystems with the expansion rates $H_{n}$. Due to the heat exchange at fixed total energy $E_{V}$ these states are organized in the equilibrium thermal state, which corresponds to the coherent de Sitter expansion of the whole system. The coherence due to thermalization may explain the horizon problem, i.e. why the causally-disconnected regions of the CMB are in thermal equilibrium.
IV.4 Regularization vs thermalization
@@
One may expect that the heat exchange between subsystems leads not only to the coherence of the de Sitter states, but to the general coherence of the metric fields, when the metric fields $g_{\mu\nu(n)}$ of the subsystems become equal, thus forming the common metric $g_{\mu\nu}$. If this is true, this could be some kind of the thermodynamic gravity, but without using the holographic principle.
IV.5 Coherence vs thermalization
@@
At first glance this formation of the time dependent coherent state looks similar to the formation of the coherent spin precession (magnon BEC) from the incoherent precessions about the local magnetic fields with random frequencies $\omega_{n}$.BunkovVolovik2013 The latter corresponds to random $H_{n}$. The coherence develops due to the spin currents between the regions of local precessions – the analog of the heat exchange. The common frequency $\omega$ of the coherent precession corresponds to the common Hubble parameter $H$. They have the same dimension, $[H]=[\omega]=[1/t]$, while the dimensionless magnon number (or the spin projection $S_{z}$) corresponds to the entropy which is also dimensionless. However, the source of the coherence is the exchange of spin instead of exchange of the energies, and also the formation of coherent state is due to minimization of the total energy $E$ at fixed projection $S_{z}$ of the total spin on magnetic field. This leads to the common frequency, which plays the role of the chemical potential $\mu$, which emerges due to the exchange of particles between the regions. That is why this process is quite opposite to the formation of temperature and entropy, where energy is fixed, while entropy reaches maximum value. This process corresponds to the minimization of energy at fixed total entropy.
Anyway, in all the cases the parameters, which are the same in all subsystems in equilibrium – common frequency $\omega$, common chemical potential $\mu$, common temperature $T$, common angular velocity $\boldsymbol{\Omega}$ and common Hubble parameter $H$ – have the same dimension, $[T]=[H]=[\mu]=[\omega]=[\Omega]=[1/t]$. Note also that together with the variables $T$, $\mu$ and $\boldsymbol{\Omega}$, the Hubble parameter $H$ enters the chiral anomaly effects, see Sec. IV.6.
IV.6 de Sitter contribution to chiral anomaly
Let us consider the chiral vortical effect – the appearance of chiral current in fermionic system in the presence of rotation. For Dirac fermions in the flat spacetime the chiral current ${\bf j}_{A}$ contains the following contributions:Vilenkin1979 ; Vilenkin1980 ; Kharzeev2016 ; Stone2018 ; Abramchuk2024
${\bf j}_{A}=\left(\frac{T^{2}}{6}+\frac{\mu^{2}}{2\pi^{2}}+\frac{\Omega^{2}}{24\pi^{2}}\right)\boldsymbol{\Omega}\,.$ (35)
In the de Sitter state the curvature term can be added.Flachi2018 ; Khakimov2024 According to Ref. Khakimov2024 the gradient expansion gives:
${\bf j}_{A}=\frac{H^{2}}{8\pi^{2}}\boldsymbol{\Omega}\,.$ (36)
The same term is obtained for the anti-de Sitter spacetime where $H^{2}<0$.Ambrus2021 In Ref.Flachi2018 the curvature term comes from the shift of the fermionic mass gap due to curvature, and this leads to the Eq.(36) with opposite sign.
Comparison of Eq.(36) with Eq.(35) demonstrates that there is the difference between the contribution $T^{2}/6$ in Eq.(35), which comes from the temperature of matter, and the contribution $T^{2}/8$ from the temperature $T=H/\pi$ of the de Sitter environment in Eq.(36). However, it is possible that the more rigorous calculations (beyond the gradient expansion, with the proper conservation laws and with the proper limit cases) can modify the coefficient in Eq.(36). It is not excluded that these two contributions may cancel each other when the matter and the gravitational background are in equilibrium and thus have the same temperature. Such cancellation of the currents generated by two different mixed gravitational anomalies in rotating chiral liquid under the full equilibrium was found in Ref. NissinenVolovik2022 . This supports the Bloch theorem on the absence of the total current in equilibrium, see also Refs. Yamamoto2015 ; Volovik2017 ; Zubkov2018 .
Anyway, the participation of the curvature ${\cal R}=-12H^{2}$ in the thermodynamics of the chiral anomaly demonstrates the uniqueness of the de Sitter state in providing the thermal-like response.
V Contracting de Sitter and black hole entropy
V.1 de Sitter vs black hole
@@ V Contracting de Sitter and black hole entropy
V.2 Entropy of expanding, contracting and static de Sitter
Let us start with the fully static de Sitter metric:
$ds^{2}=-(1-H^{2}r^{2})dt^{2}+\frac{dr^{2}}{1-H^{2}r^{2}}+r^{2}d\Omega^{2}\,.$ (37)
This metric has singularity at the cosmological horizon, which requires the proper attention. In one approach this static configuration can be considered as the intermediate state between the two opposite states – the expanding de Sitter state with $H>0$ and the contracting de Sitter state with $H<0$ (see the same consideration of the static hole as the intermediate state between black and white holes in Ref. Volovik2022 ). In the contracting de Sitter vacuum with $H<0$ the local temperature is negative, $T=H/\pi<0$, and the local entropy is also negative, $s_{\rm contracting}=3H/4G=-3/(4Gr_{h})<0$. The total entropy in the Hubble volume of contracting de Sitter is
@@ V.2 Entropy of expanding, contracting and static de Sitter
where $A$ is the area of the cosmological horizon. That is why the entropy of the static de Sitter as the intermediate state is zero, and correspondingly the temperature must be infinite, see also Ref.Milekhin2024 .
V.3 Gravastar – black hole with de Sitter core
Let as consider the black hole, which is obtained by the deformation of the gravastar.Volovik2023a The gravastar is the object, which contains the de Sitter state inside the black hole horizon.Chapline2003 ; Mazur2023 ; Mottola2023 We consider the gravastar in which the black hole horizon coincides with the de Sitter horizon, $r_{\rm bh}=r_{h}$ (the metric in the state with the critical value of the mass parameter $m=2MG|H|=1$ at which two horizons merge is illustrated in Fig. 1 of Ref. Dymnikova2002 ). Such gravastar is stationary, that is why it has no Hawking radiation and its entropy is zero. In the Painlevé-Gullstrand form the metric of such gravastar is given by Eq.(1) with the following shift velocity:
@@ V.3 Gravastar – black hole with de Sitter core
V.4 Entropy of black hole from negative entropy of contracting de Sitter
According to Eq.(38) the region of the contracting de Sitter core of the gravastar has negative entropy, $S_{\rm contracting}=-A/4G$. That is why the gravastar is unstable towards the shrinking of the de Sitter region, since this leads to the increase of the entropy of the whole system due to decrease of the negative entropy of the contracting de Sitter state. Due to the energy conservation the shrinking of the volume of the de Sitter core leads to formation of the singularity ar $r=0$, where the mass becomes concentrated. In the final state – the black hole – the de Sitter region with negative entropy fully disappears by shrinking to the singularity with mass $M$. The resulting black hole with mass $M$ acquires the positive entropy, $A/4G$:
@@ V.4 Entropy of black hole from negative entropy of contracting de Sitter
The black hole entropy is also concentrated in the singularity, together with the curvature ${\cal R}$ and mass $M$.Volovik2023a This again supports the holographic connection between the entropy of bulk, which is concentrated in the singularity, and the surface entropy of the black hole horizon.
Eq.(41) allows to interpret the zero value of the entropy of the gravastar in terms of the cancellation of entropies of two horizons. In the initial gravastar state the entropy of the contracting de Sitter horizon fully compensates the entropy of the black hole horizon:
@@
V.5 White hole and anti-gravastar
@@
It is obtained from the pure white hole, which has the negative horizon entropy $S_{\rm WH}=-A/4G$,Volovik2022 ; Volovik2020 by growing the de Sitter state in its core with positive entropy. In the anti-gravastar state the entropies of two horizons cancel each other in the same way as in the gravastar:
@@
It is important that the singularity in the metric of gravastar and in the metric of anti-gravastar can be smoothly removed by deformation. That is why these objects do not depend on the choice of the coordinate systems, and thus they are equivalent to the fully static black hole with the fully static de Sitter core:
$ds^{2}=-(1-2GM/r)dt^{2}+\frac{dr^{2}}{1-2GM/r}+r^{2}d\Omega^{2}\,\,,\,\,r>2MG=1/H\,,$ (46)
$ds^{2}=-(1-H^{2}r^{2})dt^{2}+\frac{dr^{2}}{1-H^{2}r^{2}}+r^{2}d\Omega^{2}\,\,,\,\,r<2MG=1/H\,.$ (47)
V.6 Heat exchange between black holes in the multi-metric ensemble
@@
VI Thermodynamics of de Sitter state and $f(R)$ gravity
VI.1 Thermodynamic variables in $f(R)$ gravity
@@
VI.2 Gibbs-Duhem relation in $f(R)$ gravity
@@
VI.3 Entropy of cosmological horizon in terms of effective gravitational coupling
@@
VI.4 Example of quadratic gravity
@@
VII From de Sitter thermodynamics to de Sitter decay
VII.1 de Sitter state as thermal bath for matter
@@ VII From de Sitter thermodynamics to de Sitter decay
The nonzero local temperature of the gravitational vacuum suggests that the de Sitter vacuum is locally unstable towards the creation of thermal matter from the vacuum by thermal activation, see Section II.2. This is distinct from the mechanism of creation of the pairs of particles by Hawking radiation from the cosmological horizon, which may or may not lead to the decay of the vacuum energy. There are still controversies concerning the stability of the de Sitter vacuum caused by Hawking radiation, see e.g. Kamenshchik2022 ; Polyakov2012 ; Polyakov2022 and references therein.
VII.2 de Sitter decay due to thermalization of matter by de Sitter heat bath
To describe the decay of the vacuum due to activation and thermalization of matter, the extension of the Starobinsky analysis of the vacuum decayStarobinsky1994 ; Starobinsky1996 ; Starobinsky1997 ; Starobinsky2023 is needed. Especially the revolutionary stochastic inflation approach pioneered by StarobinskyStarobinsky1982 ; Starobinsky1986 is extremely useful, although it requires some modifications.Pattison2019 ; Cruces2022 ; Cruces2022b ; Shellard2024 This also includes the so-called separate universe approach, which is somewhat similar to the multi-metric gravity discussed in Sec. IV. Here we consider the simple phenomenological scenario based on the energy exchange between the thermal de Sitter and the thermal matter.
@@
Here we use for simplicity the conventional general relativity with $K=K_{0}$ and $\epsilon_{\rm vac}=\Lambda=-\frac{1}{2}K{\cal R}=6KH^{2}$.
This phenomenological description of the energy exchange between the de Sitter vacuum and matter does not depend on the details of the microscopic (UV) theory, and requires only the slow-roll condition, i.e. the slow variation of the Hubble parameter, $|\dot{H}|\ll H^{2}$, or which is the same $|\dot{T}|\ll T^{2}$.
Since the vacuum energy density is $\epsilon_{\rm vac}=6KH^{2}$, one obtains from Eq.(65) the following time dependence of the Hubble parameter and of energy densities:
$H(t)=b^{-1/3}M_{\rm Pl}\left(\frac{t_{\rm Pl}}{t+t_{0}}\right)^{1/3}\,,$ (66)
$\epsilon_{\rm vac}(t)=6b^{-2/3}M_{\rm Pl}^{4}\left(\frac{t_{\rm Pl}}{t+t_{0}}\right)^{2/3}\,,$ (67)
$\epsilon_{M}(t)=bH^{4}=b^{-1/3}M_{\rm Pl}^{4}\left(\frac{t_{\rm Pl}}{t+t_{0}}\right)^{4/3}\,.$ (68)
@@
$H(t=0)=b^{-1/3}M_{\rm Pl}\left(\frac{t_{\rm Pl}}{t_{0}}\right)^{1/3}\ll M_{\rm Pl}\,.$ (69)
@@
So, the scenario of thermalization of matter by the de Sitter heat bath gives Eqs. (64) and (65). They produce the inflation in terms of two phenomenological parameters, $b$ and $t_{0}$, which determine the decay of the vacuum energy density in Eq.(67).
VII.3 Connection to holographic principle
The evolution of the gravitational (dark) energy in Eq. (67) and of the energy of the relativistic matter in Eq.(68) allows us to consider another proposal made by Padmanabhan Padmanabhan2012 (see also Refs.Krishna2022 ; Prasanthan2024 and references therein). It is the holographic postulate which connects the expansion of the Hubble volume with the difference between the number $N_{\rm hor}$ of microstates on the surface of horizon (one degree of freedom per Planck area) and the number $N_{\rm bulk}$ of the degrees of freedom in bulk:
$\frac{dV_{H}}{dt}=G\,(N_{\rm hor}-N_{\rm bulk})\,.$ (70)
This postulate suggests that the expansion of the universe is being driven towards the holographic equipartition, so that in the equilibrium de Sitter state all the bulk degrees of freedom inside the horizon can be expressed via the horizon degrees of freedom, $N_{\rm bulk}=N_{\rm hor}$.
Instead of the speculative degrees of freedom, we consider here the corresponding thermodynamic entropy. We already found that according to Eq.(18) the gravitational entropy of the Hubble volume is equal to the entropy related to horizon, $S_{\rm bulk}=S_{\rm hor}$. Let us consider what happens in the presence of matter. On one hand, matter adds its contribution $S_{M}$ to the bulk entropy, and thus $S_{\rm bulk}=S_{M}+S_{\rm hor}$. On the other hand, matter violates the de Sitter symmetry, which leads to the time dependence of the Hubble volume.
The entropy density of matter $s_{M}(t)$ can be obtained from the matter energy density in Eq.(68):
$s_{M}(t)=\frac{4\pi}{3}bH^{3}(t)\,.$ (71)
Then the total entropy of matter $S_{M}$ in the Hubble volume is time independent:
$S_{M}=s_{M}(t)V_{H}(t)=\left(\frac{4\pi}{3}\right)^{2}b\,.$ (72)
The time derivative of the Hubble volume $dV_{H}/dt$, which is obtained from Eq.(66), is also time independent:
$\frac{dV_{H}}{dt}=\frac{4\pi}{3}bM_{\rm Pl}^{2}\,.$ (73)
Then comparing Eq.(73) with Eq.(72) and using equation $S_{M}=S_{\rm bulk}-S_{\rm hor}$ one obtains the general relation between the expansion of the Hubble volume and the difference between the bulk entropy and and its holographic surface value:
$\frac{dV_{H}}{dt}=\frac{3}{4\pi}l_{\rm Pl}^{2}S_{M}=12G\,(S_{\rm bulk}-S_{\rm hor})\,.$ (74)
Up to the sign and numerical coefficient, this coincides with Eq.(70), which supports the Padmanabhan holographic conjecture.Padmanabhan2012
VII.4 de Sitter decay and Zel’dovich stiff matter
As was mentioned by Padmanabhan,Padmanabhan2005 his model leads to a late time cosmological constant in Eq.(67), which is independent of the initial value, but its value is still far too large. Can we fix this? In Sec. VI.2 we obtained that the thermodynamics of de Sitter thermal bath corresponds to the Zel’dovich stiff matter with $w=1$. Let us try the stiff matter scenario using the phenomenological approach.